OpenStudy (anonymous):

Two consecutive even integers have a sum of 38 . Find the integers.

OpenStudy (welshfella):

if one number is x then next one is x +1 so we can form the equation x + x + 1 = 38 solve for x

OpenStudy (welshfella):

oops sorry I misread the question as they are both even numbers that will be x and x + 2 so x + x + 2 = 38

OpenStudy (anonymous):

x+ (x+1) = 38 2x+1 = 38 2x=37 x= 18.5

OpenStudy (anonymous):

oh oops hold on

OpenStudy (welshfella):

- you made same mistake as me!

OpenStudy (anonymous):

x=18

OpenStudy (anonymous):

yeahh haha =)

OpenStudy (welshfella):

right so the other one is ?

OpenStudy (anonymous):

I get stuck after I find x lol the explanation froze up and that's as far as told me to do =o

OpenStudy (welshfella):

one number is x = 18 the next even number = 18 + 2

OpenStudy (anonymous):

oh ok that's easy so my numbers are 18,20

OpenStudy (anonymous):

thank you! @welshfella

OpenStudy (welshfella):

yes yw